OpenStudy (anonymous):

How many 1/12s are in 2?

OpenStudy (mkmkasim):

2

OpenStudy (mkmkasim):

24 sorry

OpenStudy (anonymous):

\[ \frac {1}{12}\cdot x = 2 \]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Do u understand why its 24, it takes 12/12= 1 to equal one, so it takes 24 twelves to equal two :) Using an equation like @wio is always good to solve an unknown variable :)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

These are the steps to solve Divide 1/12 from both sides (1/12)/x= 2/(1/12) Then multiply x= 2*12 x= 24 :)

OpenStudy (anonymous):

Any other q's?

OpenStudy (anonymous):

What is 9 divided by 2/3

OpenStudy (anonymous):

(9) / (2/3) Flip the reciprocal 9 x 3/2 Multiply 27/2 or 13.5

OpenStudy (anonymous):

Understand?